# Ranked from Within: Ranking Large Multimodal Models for Visual Question Answering Without Labels

### v1

@@ Abstract
As large multimodal models (LMMs) are increasingly deployed across diverse applications, the need for adaptable, real-world model ranking has become paramount. Traditional evaluation methods are largely dataset-centric, relying on fixed, labeled datasets and supervised metrics, which are resource-intensive and may lack generalizability to novel scenarios, highlighting the importance of unsupervised ranking. In this work, we explore unsupervised model ranking for LMMs by leveraging their uncertainty signals, such as softmax probabilities. We evaluate state-of-the-art LMMs (e.g., LLaVA) across visual question answering benchmarks, analyzing how uncertainty-based metrics can reflect model performance. Our findings show that uncertainty scores derived from softmax distributions provide a robust, consistent basis for ranking models across varied tasks. This finding enables the ranking of LMMs on real-world, unlabeled data for visual question answering, providing a practical approach for selecting models across diverse domains without requiring manual annotation.
@@ 1 Introduction
Large Multimodal Models (LMMs), such as LLaVA [46] and InstructBLIP [11], have demonstrated remarkable capabilities in handling a wide array of complex multimodal tasks, proving highly adaptable across diverse real-world applications [44, 6, 11, 8, 5, 31]. From addressing scientific challenges [47, 22] and performing optical character recognition [55, 52, 53] to identifying spatial positions of objects [70, 77], LMMs are increasingly essential in practical settings. As LMMs evolve, the need for rigorous evaluation metrics that accurately capture their capabilities and limitations has become more urgent. Thus, numerous benchmarks have been developed [47, 22, 52, 67, 53, 77, 80], aiming to provide reliable rankings and guide users in selecting models best suited for specific deployment scenarios.
@@
However, these benchmarks rely on carefully curated datasets, which can require substantial resources to develop and label. For many users who may not have access to these resources, assessing model performance can be challenging. Additionally, standard evaluations are often dataset-centric, depending on fixed, human-labeled metrics that may not capture the full range of model capabilities necessary for diverse applications. As LMM tasks continue to diversify and expand, evaluating them effectively has become increasingly complex, with new applications frequently needing additional data curation and specialized capabilities.
Addressing this challenge is essential for developing adaptable and scalable evaluation methods that support the safe and effective deployment of LMMs across diverse real-world environments. As shown in Fig. 1, selecting the most suitable LMM from a range of available options is challenging in the absence of annotations. To tackle this, our work proposes a proxy ranking metric that closely aligns with true performance, enabling effective model comparison in label-scarce settings. In image classification, several studies have aimed to predict a model’s performance on out-of-distribution (OOD) datasets using in-distribution (ID) metrics [68, 54, 4]. Known as accuracy-on-the-line (AoL), this approach demonstrates a strong linear correlation between probit-scaled ID accuracy and OOD accuracy across different classifiers. For ranking LMMs, we can leverage the AoL phenomenon [54] as a starting point. However, defining ID performance for LMMs is often challenging, as training datasets may be proprietary and training tasks may differ from test tasks. Therefore, we utilize performance on existing benchmarks as a practical basis for ranking models in new testing domains.
@@
To explore this further, this paper begins by conducting a correlation study between performance on different benchmarks to examine the effectiveness of AoL for ranking models. With the observed instability and inconsistency of AoL for ranking LMMs, we then investigate the effectiveness of prediction uncertainty as the ranking indicator. LMMs generate answers in an open-form by producing token sequences, where each token is sampled from a probability distribution over the vocabulary. This enables assessing model uncertainty using logits on different token positions. To this end, we evaluate LMMs from different series. They have different training frameworks, such as LLaVA-V1.5 [44] and InstructBLIP [11], different visual encoders (e.g., CLIP [61] and SigLIP [81]) and language models (e.g., Vicuna [69] and LLaMA [71]). We investigate three perspectives to explore model uncertainty, which are softmax probability, self-consistency and labeled proxy set. We evaluate the ranking performance on widely-adopted multimodal benchmarks, which span in diverse domains, including reasoning scientific questions, recognizing optical characters and identifying objects’ spatial positions. Our main observations are summarized as below:
@@
We demonstrate that model performance on one dataset may not accurately reflect the model rankings on a different dataset (Section 4).
We evaluate various ranking methods and observe that their effectiveness is influenced by task characteristics (i.e., closed-form or free-form generation). Among the methods considered, NLLmin stands out as the most effective (Section 5).
@@
When examining correlations in model performance across different dataset pairs, we observe that variations in text prompt similarity align more closely with performance correlation changes than variations in image feature similarity (Section 6).
@@ 2 Related Work
The goal of this task is to rank and select a best performant models without the access to the data annotations of target environments. The research on this task can date back to Forster et al. [14], and was further investigated in various tasks: (1) outlier detection [84, 83]; (2) image classification [36, 73, 85, 4, 54, 64]; (3) time series anomaly detection [78]; (4) multivariate anomaly detection in manufacturing systems [13], etc.
We discuss the most relevant two lines of research as follows. Miller et al. [54] introduce an accuracy-on-the-line (AoL) phenomenon where strong linear correlation between probit-scaled in-distribution (ID) accuracy and out-of-distribution (OOD) accuracy across a variety of ML models. This means ID accuracy serves as a good indicator of model performance for a target domain. Shi et al. [64] revisit the established concept of lowest common ancestor (LCA) distance, which measures the hierarchical distance between labels and predictions within a predefined class hierarchy. By the observed linear correlation between ID LCA distance and OOD accuracy, it is viable to select a model based LCA distance of predictions. This work differs from prior study that we focus on ranking LMMs. This group of models makes predictions in a significantly different way from conventional classification models, which also results in different evaluation protocols.
@@
Uncertainty estimation aims to measure an ML model’s confidence level of a certain prediction [20]. There are two major types of uncertainty one can model: aleatoric and epistemic uncertainty [33]. Aleatoric uncertainty refers to the noise inherent in the observations; while epistemic uncertainty accounts for uncertainty in the model parameters. To quantify uncertainty, uncertainty estimation are roughly categorized into four groups: (1) single deterministic method [20, 21]; (2) model ensembling [39]; (3) Bayesian method [16, 10, 38]; (4) test-time augmentation [48, 75]. Single deterministic methods calculate prediction uncertainty based on one forward pass within a deterministic ML model. By contrast, Bayesian methods only leverage the inherent stochasticity of an ML model (e.g.,, dropout layer in deep neural networks [16]). We primarily use (1) and (3) for ranking LMMs due to their simplicity in implementation and model- and domain-agnostic.
@@
The rapid development of LMMs has greatly propelled advancements in multimodal models, showcasing significant improvements in their perception and reasoning abilities. This shift has rendered traditional benchmarks, which focus solely on isolated task performance [32, 3], inadequate for thorough evaluation. Researchers have introduced new benchmarks aimed at evaluating LMMs across a broad spectrum of multimodal tasks [19, 43, 63]. However, many benchmarks lack detailed assessments and robust evaluation measures. Recent studies [80, 77] highlight the need for more comprehensive benchmarks to evaluate the reasoning and understanding abilities of LMMs effectively.
@@ 3 Task Formulation
Let a multi-modal task be represented by a dataset , where and denote the input prompt and the corresponding answer for the -th sample, respectively. We have access to large multimodal models (LMMs), denoted as . Each LMM , with pre-trained weights , generates a sequence of tokens from the input prompt via a decoding process: , where represents the -th generated token. To assess the performance of each LMM, this task employs an evaluation metric (e.g., accuracy) that determines the ground-truth performance by comparing the generated sequences with the ground-truth answers .
The objective of this paper is to develop methods for computing scores for each LMM without requiring access to task-specific data annotations. Ideally, these computed scores should closely correlate with ground-truth performance, allowing us to rank and select LMMs based on their performance using only these scores.
@@
We use Spearman’s rank correlation coefficient () [34] to assess the monotonic relationship between the computed scores and model performance. Additionally, we calculate the weighted version of Kendall’s rank correlation () [66], which has been shown to effectively identify top-ranked items of interest [79]. Both coefficients range from , where values close to or indicate strong negative or positive correlations, respectively, and a value of denotes no correlation.
@@ 4 Uncertainty for Ranking LMMs
In this section, we discuss the unique characteristics of ranking various Large Multimodal Models (LMMs) compared to conventional ML models. We then present three distinct approaches that leverage uncertainty in model predictions for ranking.
@@ 4.1 What makes ranking LMMs interesting?
Unique Challenges and Potential Methods for Ranking LMMs.
While LMMs can be considered a subset of machine learning (ML) models, ranking them introduces unique challenges not present in traditional ML models. Below, we discuss the unique characteristics of LMMs and the challenges they bring to the model-ranking process.
Scale and Architectural Complexity. LMMs, often pre-trained with billions of parameters (e.g., LLaVA-V1.5-7B [46] with billion parameters), present significant challenges for traditional “white-box” analyses. Techniques such as examining neuron activations [2], analyzing attention patterns [17, 74], or assessing weight norms across layers [57] are computationally and manually prohibitive due to their sheer scale.
@@
Limited Transparency. LMMs also exhibit transparency challenges. First, these models are typically trained on large datasets of instruction fine-tuning samples, which may be publicly accessible or proprietary data. As a result, risk assessment methods that require access to training data (e.g., uncertainty estimation [20]) are either unsuitable or should be adjusted. Additionally, although some LMMs may disclose their final model weights, other information, such as training loss and intermediate checkpoints, often remains undisclosed. The lack of the access to training data and training details limits the use of techniques in the unsupervised accuracy estimation [12, 72] and scaling law [28, 29].
Using Accuracy-on-the-Line (AoL)?
Given these challenges, one may consider leveraging the AoL phenomenon as a potential method for ranking LMMs. Miller et al. [54] describe AoL as a strong linear correlation between probit-scaled in-distribution (ID) accuracy and out-of-distribution (OOD) accuracy across various ML models. This suggests that ID accuracy could serve as a reliable predictor of OOD performance, making AoL an intuitive starting point for selecting LMMs in target testing environments.
@@
However, we suggest that AoL may not be suitable in the context of ranking LMMs for real-world deployment. First, obtaining ID performance data is often impractical, as LMMs are typically trained on large, sometimes proprietary datasets, limiting direct access to ID metrics. Second, while performance on existing benchmarks is more readily available, using this data to rank models for new deployment environments poses significant challenges. Figure 2 illustrates the correlation between proxy benchmark performance and target testing environments, revealing substantial variability in correlation strength, ranging from as high as to a negative correlation of . This inconsistency highlights the unreliability of using benchmark performance as a sole ranking criterion. Third, relying solely on proxy benchmark performance fails to capture the unique statistical characteristics of target testing datasets. AoL tends to rank models identically across different environments, regardless of the actual deployment context, making it less suitable for selecting LMMs intended for diverse, fine-tuning-free environments.
Last, a practical alternative could involve annotating a subset of the target environment and using the performance on this labeled subset as a ranking indicator, if the budget allows. This method might be useful in scenarios with non-stationary distributions, as periodic re-annotation may be needed to reflect changes in the testing environment. However, detecting such shifts requires carefully designed methods, and in specialized domains like medical image analysis [62] or harmful content detection [35], annotation can be both time-consuming and require expert knowledge.
What Are the Possible Methods?
As model size and architectural complexity increase, traditional methods become less feasible due to the substantial computational demands and the need for more sophisticated algorithm design. Techniques that leverage scaling laws [30] also require multiple models from the same family, but with varying sizes, to train a regressor that estimates model performance. In model ranking contexts, computational overhead grows substantially with the number of models, making such methods impractical for large-scale deployment.
In contrast, model outputs—such as the logits used for token prediction and the generated tokens—are readily accessible without requiring deep examination of the model architecture or complex information extraction techniques. Therefore, we focus primarily on methods that utilize the model’s prediction logits and output strings to rank LMMs.
4.2 Uncertainty by Softmax Probability
@@
Unlike models used in image classification, where predictions correspond to the entry with the highest softmax probability, LMMs generate answers in an open-form manner by producing sequences of tokens. However, each token generation can still be viewed as a classification process, where the model selects the token with the highest probability over its entire vocabulary. This allows us to assess model uncertainty based on the logits at each token position.
@@
For token-level uncertainty, we can focus on two specific positions: the first token and the penultimate token (i.e., the token preceding the end-of-sequence token), as illustrated in Figure 3. The logit of the first token reflects the model’s initial response to the input prompt, while the logit of the penultimate token captures the model’s understanding of both the prompt and the generated response. The uncertainty associated with these two positions may provide insight into the model’s overall confidence in answering the question. Beyond token-level uncertainty, sentence-level uncertainty can be calculated by aggregating uncertainty across all tokens in the sequence, as suggested in [50, 23]. Specifically, sentence-level uncertainty can be quantified using the mean or the maximal negative log likelihood values across the entire generated sequence:
@@
where is the likelihood of the word generated by the LMM at the -th token of the -th sentence and is the number of tokens generated for this sentence. Equation 1 quantifies a sentence’s uncertainty by identifying its least likely token, while the average per-token uncertainty (Equation 2) normalizes the log probability across sequences of varying lengths, allowing for length-independent comparisons of uncertainty [49, 56]. In addition, Equation 2 is closely related to [25, 51], which provides a comparable measure of model quality.
@@
Alternatively, entropy can also be used instead of negative log likelihood to assess uncertainty. In the context of model ranking, we adopt normalized entropy to account for varying vocabulary sizes across models, thus scaling entropy to the interval and making it comparable across different models. The normalized entropy is given by:
@@
There are eight variants of output probability-based methods, denoted as NLLF, NLLP, NLLmin, NLLavg, EntF, EntP, Entmax, and Entavg. “NLL” and “Ent” represent negative log likelihood and entropy, respectively, while “F” and “P” refer to the first and penultimate tokens.
4.3 Uncertainty by Self-Consistency

### v2

# Ranked from Within: Ranking Large Multimodal Models Without Labels

@@ Abstract
Can the relative performance of a pre-trained large multimodal model (LMM) be predicted without access to labels? As LMMs proliferate, it becomes increasingly important to develop efficient ways to choose between them when faced with new data or tasks. The usual approach does the equivalent of giving the models an exam and marking them. We opt to avoid marking and the associated labor of determining the ground-truth answers. Instead, we explore other signals elicited and ascertain how well the models know their own limits, evaluating the effectiveness of these signals at unsupervised model ranking. We evaluate state-of-the-art LMMs (e.g., LLaVA) across visual question answering benchmarks, analyzing how well uncertainty-based metrics can predict relative model performance. Our findings show that uncertainty scores derived from softmax distributions provide a robust and consistent basis for ranking models across various tasks. This facilitates the ranking of LMMs on unlabeled data, providing a practical approach for selecting models for diverse target domains without requiring manual annotation.
@@ 1 Introduction
Large multimodal models (LMMs), such as LLaVA (Liu et al., 2024c) and InstructBLIP (Dai et al., 2023), have demonstrated remarkable capabilities in handling a wide array of complex multimodal tasks, proving highly adaptable across diverse real-world applications (Liu et al., 2024a; Dai et al., 2023; Wang et al., 2024a; Huang et al., 2024, 2025; Zheng et al., 2025). From addressing scientific challenges (Lu et al., 2022; Hiippala et al., 2021) and performing optical character recognition (Mishra et al., 2019; Masry et al., 2022; Mathew et al., 2021) to identifying the spatial position of objects (Tong et al., 2024; x.ai, 2024), LMMs are increasingly widespread in practical settings. As LMMs proliferate, the need for rigorous evaluation metrics that accurately capture their capabilities and limitations has become more urgent. Thus, numerous benchmarks have been developed (Lu et al., 2022; Hiippala et al., 2021; Masry et al., 2022; Singh et al., 2019; Mathew et al., 2021; x.ai, 2024; Yue et al., 2024), aiming to provide reliable rankings and guide users in selecting models best suited for specific deployment scenarios.
@@
However, these benchmarks are based on carefully curated datasets that can require substantial resources to develop and label. For many users who may not have access to these resources, assessing model performance can be challenging. Additionally, standard evaluations are often dataset-centric, depending on fixed, human-labeled metrics that may not capture the full range of model capabilities necessary for diverse applications. As LMM tasks continue to diversify and expand, evaluating them effectively has become increasingly complex, with new applications frequently needing additional data curation and specialized capabilities.
Addressing the challenge of efficiently evaluating a set of LMMs is important for their usability and effectiveness in deployment. As illustrated in Figure 1, selecting the most suitable LMM from a range of available options is challenging in the absence of annotations. To address this, our work investigates label-free proxy ranking scores that closely align with the true performance ranking, enabling effective model comparison in label-scarce settings.
Our first finding is that the naïve approach of using the model’s performance on an existing benchmark to rank its performance in a new target environment could be unstable. We further report that measures of prediction uncertainty are more effective ranking indicators. LMMs generate answers in an open-ended form by producing token sequences, with each token selected from the vocabulary based on the probability. This enables model uncertainty to be assessed using the logits at each token position.
@@
To this end, we evaluate different LMMs that have different training frameworks, e.g., LLaVA-V1.5 (Liu et al., 2024a) and InstructBLIP (Dai et al., 2023), different visual encoders, e.g., CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023), and language models, e.g., Vicuna (Team, 2023) and LLaMA (Touvron et al., 2023). We investigate three categories of model uncertainty approaches: softmax probabilities, self-consistency, and labeled proxy sets. We evaluate the ranking performance on widely-adopted multimodal benchmarks, which span diverse domains, including reasoning scientific questions, recognizing optical characters, and identifying objects’ spatial positions. Our main findings are that
@@
the performance of models on one dataset may not accurately reflect the ranking of the same models on a different dataset (Section 4);
the effectiveness of ranking methods is influenced by task characteristics (e.g., closed-form or free-form generation), but probability-based variants are typically quite robust and predictive (Section 5); and
@@
when examining correlations in model performance across different dataset pairs, we observe that text prompt similarity better correlates with model performance across datasets than image feature similarity (Section 6).
@@ 2 Related Work
The goal of this task is to rank and select a best performant models without the access to the data annotations of target environments. The research on this task can date back to Forster et al. (Forster, 2000), and was further investigated in various tasks: (1) outlier detection (Zhao et al., 2022, 2021); (2) image classification (Kotary et al., 2022; Tu et al., 2024a; Zohar et al., 2024; Baek et al., 2022; Miller et al., 2021; Shi et al., 2024a); (3) time series anomaly detection (Ying et al., 2020); (4) multivariate anomaly detection in manufacturing systems (Engbers & Freitag, 2024), etc.
We discuss the most relevant two lines of research as follows. Miller et al. (2021) introduce an accuracy-on-the-line (AoL) phenomenon where strong linear correlation between probit-scaled in-distribution (ID) accuracy and out-of-distribution (OOD) accuracy across a variety of ML models. This means ID accuracy serves as a good indicator of model performance for a target domain. Shi et al. (2024a) revisit the established concept of lowest common ancestor (LCA) distance, which measures the hierarchical distance between labels and predictions within a predefined class hierarchy. By the observed linear correlation between ID LCA distance and OOD accuracy, it is viable to select a model based LCA distance of predictions. This work differs from prior study that we focus on ranking LMMs. This group of models makes predictions in a significantly different way from conventional classification models, which also results in different evaluation protocols.
@@
Uncertainty estimation seeks to quantify an ML model’s confidence in its predictions (Guo et al., 2017). Recent studies have been dedicated to exploring uncertainty estimation specifically for LLMs (Kuhn et al., 2023; Malinin & Gales, 2021; Xiao & Wang, 2021; Huang et al., 2023; Lin et al., 2023; Kadavath et al., 2022; Azaria & Mitchell, 2023; Gottesman & Geva, 2024). For instance, Xiao et al. (2021) utilize ensemble methods to evaluate uncertainty in natural language generation models. Malinin et al. (2021) similarly introduce a unified approach to uncertainty estimation, leveraging ensemble methods, for autoregressive structured prediction tasks. To deal with the challenge of capturing “semantic equivalence” in natural language, Kuhn et al. (2023) propose semantic entropy, a method that utilizes linguistic invariances derived from shared meanings. Additionally, internal states of LMMs can be leveraged for uncertainty quantification or error detection by training a classifier (Azaria & Mitchell, 2023; Gottesman & Geva, 2024). This paper does not propose a new way to estimate uncertainty. Instead, it offers the novel insight that the existing uncertainty in LMM-generated outputs effectively reflects their relative performance across benchmarks without manual labels.
@@
The rapid development of LMMs has greatly propelled advancements in multimodal models, showcasing significant improvements in their perception and reasoning abilities. This shift has rendered traditional benchmarks, which focus solely on isolated task performance (Karpathy & Fei-Fei, 2015; Antol et al., 2015). Researchers have introduced new benchmarks to evaluate LMM in a broad spectrum of multimodal tasks (Goyal et al., 2017; Lin et al., 2014; Russakovsky et al., 2015). Recent studies (Yue et al., 2024; x.ai, 2024; Fu et al., 2023) highlight the need for more comprehensive benchmarks to effectively evaluate the reasoning and understanding abilities of LMMs. For instance,
Several benchmarks (x.ai, 2024; Ainslie et al., 2023; Tong et al., 2024) have been developed to assess multimodal models’ real-world spatial understanding. Lu et al. (2024) and Zhang et al. (2024) introduce benchmarks specifically designed to evaluate MLLMs’ mathematical reasoning, focusing on their ability to comprehend and reason about visual mathematical figures. Yue et al. (2024) carefully curate a diverse set of multi-discipline tasks that require college-level subject knowledge and complex reasoning. Additionally, numerous benchmarks (Mishra et al., 2019; Masry et al., 2022; Mathew et al., 2021; Liu et al., 2023) assess the performance of LMMs in optical character recognition.
@@ 3 Task Formulation
Let a multimodal task be represented by a dataset , where and denote the input prompt and the corresponding answer for the -th sample, respectively. We have access to large multimodal models (LMMs), denoted as . Each LMM , with pre-trained weights , generates a sequence of tokens from the input prompt via a decoding process: , where represents the -th generated token. To assess the performance of each LMM, this task employs an evaluation metric (e.g., accuracy) that determines the ground-truth performance by comparing the generated sequences with the ground-truth answers .
The objective of this paper is to develop methods for computing a score for each LMM without requiring access to task-specific data annotations. Ideally, these computed scores should closely correlate with ground-truth performance, allowing us to rank and select LMMs based on their performance using only these scores.
@@
We use Spearman’s rank correlation coefficient (Kendall, 1948) to evaluate the monotonic relationship between scores and model performance. Additionally, we calculate Kendall’s weighted rank correlation (Shieh, 1998), which effectively highlights top-ranked items (You et al., 2021). Both coefficients range from , with values near or indicating strong negative or positive correlations, and indicating no correlation.
@@ 4 Uncertainty for Ranking LMMs
This section discusses the unique characteristics of ranking various LMMs compared to conventional ML models. We then introduce three distinct approaches that leverage uncertainty in model predictions for ranking.
@@ 4.1 What Makes Ranking LMMs Interesting?
Unique Challenges for Ranking LMMs.
@@
While LMMs can be considered a subset of machine learning (ML) models, ranking them introduces unique challenges not present in traditional ML models. Below, we discuss the unique characteristics of LMMs and the challenges they bring to the model-ranking process. First, LMMs often have billions of pre-trained parameters, which presents significant challenges for traditional “white-box” analysis for ranking various LMMs. Second, these models are typically trained on large datasets of instruction fine-tuning samples, which may be proprietary data. As a result, risk assessment methods that require access to training data are either unsuitable or should be adjusted. Additionally, although some LMMs may disclose their final model weights, other information, such as training loss and intermediate checkpoints, often remains undisclosed. The lack of the access to training details limits the use of techniques in unsupervised accuracy estimation (Deng & Zheng, 2021; Tu et al., 2023).
Does Accuracy-on-the-Line (AoL) Suffice?
@@
Given these challenges, one may consider leveraging the AoL phenomenon (Miller et al., 2021) as a potential method for ranking LMMs. AoL refers to the strong linear correlation between probit-scaled in-distribution (ID) accuracy and out-of-distribution (OOD) accuracy across various ML models. This suggests that ID accuracy could serve as a reliable predictor of OOD performance, making AoL suitable for selecting LMMs in target testing environments. However, there are several reasons why this is not the case. First, obtaining ID performance data is often impractical, as LMMs are typically trained on large, sometimes proprietary datasets, limiting direct access to ID metrics. Second, while performance on existing benchmarks is more readily available, using this data to rank models for new deployment environments is problematic. Figure 2 illustrates the correlation between proxy benchmark performance and target testing environments, revealing extreme variability in correlation strength. This highlights the unreliability of using benchmark performance as the sole ranking criterion. Third, relying solely on proxy benchmark performance fails to capture the unique statistical characteristics of target testing datasets. AoL tends to rank models identically across different environments, regardless of the actual deployment context.
What can we use for ranking LMMs?
Our approach leverages the readily available outputs of LMMs—specifically, token prediction logits and generated tokens—without requiring intricate architectural analysis or complex extraction techniques. Inspired by recent work on uncertainty scores for classifier ranking (Hu et al., 2024; Tu et al., 2024b), we introduce a novel adaptation of these techniques tailored to the specific characteristics of LMMs. By analyzing the distribution of prediction logits and the variability in generated outputs, we aim to assess each model’s self-awareness of its limitations. This enables an unsupervised model ranking method, focusing primarily on techniques that utilize these readily accessible LMM outputs.
4.2 Assessing Uncertainty by Softmax Probabilities
@@
LMMs generate answers in an open-form manner by producing sequences of tokens. The selection of each token can be viewed as a classification process, where the model selects the token with, e.g., the highest probability over its entire vocabulary. This allows us to assess model uncertainty based on the logits at each token position.
@@
For token-level uncertainty (Huang et al., 2023), we can focus on two specific positions: the first token and the penultimate token (i.e., the token preceding the end-of-sequence token), as illustrated in Figure 3. The logit of the first token reflects the model’s initial response to the input prompt, while the logit of the penultimate token captures the model’s understanding of both the prompt and the generated response. The uncertainty associated with these two positions may provide insight into the model’s overall confidence in answering the question. Beyond token-level uncertainty, sentence-level uncertainty can be calculated by aggregating uncertainty across all tokens in the sequence (Manakul et al., 2023; Huang et al., 2023). Specifically, sentence-level uncertainty can be quantified using the mean or the maximum negative log-likelihood (NLL) values across the entire generated sequence:
| NLL | (1) | |||
| NLL | (2) |
@@
where is the likelihood of the word generated by the LMM at the -th token of the -th sentence and is the number of tokens generated in the sentence. Equation 1 quantifies a sentence’s uncertainty via the least likely token, while Equation 2 uses the average per-token log-likelihood, allowing for length-independent comparisons of uncertainty (Malinin & Gales, 2020; Murray & Chiang, 2018). Moreover, Equation 2 relates to perplexity, (Jelinek et al., 1977; Manning & Schütze, 1999), a standard measure of model quality.
@@
Alternatively, entropy can be used instead of the negative log-likelihood to assess uncertainty. In the context of LMMs ranking, we adopt normalized entropy to account for varying vocabulary sizes across models, thus scaling entropy to the interval and making it comparable across different models. The normalized entropy is given by:
@@
There are eight variants of output probability-based methods, denoted as NLL, NLL, NLL, NLL, Ent, Ent, Ent, and Ent. “NLL” and “Ent” represent negative log-likelihood and entropy, respectively, while “F” and “P” refer to the first and penultimate tokens.
4.3 Assessing Uncertainty by Self-Consistency